OpenStudy (cmtboy2):

Alan can chop 3/4 of a carrot in 1/8 min. How many carrots can he chop in 1 min?

OpenStudy (cmtboy2):

@EKKERKING

OpenStudy (anonymous):

@ChantySquirrel1129** @YourFavoriteRabbit

OpenStudy (anonymous):

i dont know this one

OpenStudy (hanamalik):

No clue

OpenStudy (cmtboy2):

@sky060902

OpenStudy (anonymous):

In 1 minute, he can chop (3/4) / (1/8) = (3/4)(8/1) = 6 carrots

OpenStudy (cmtboy2):

0 right

OpenStudy (anonymous):

Hmmmmm that one....Idk the answer sorry maybe you could ask some one else sorry

OpenStudy (cmtboy2):

@Diana.xL

OpenStudy (cmtboy2):